[Brahms-l] Shift report 20050201 08:00-16:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Tue Feb 01 2005 - 16:15:37 EST
No beam the entire day.  The cause of the power dip was a downed power line.  Estimate for restoration of the power line is 16:00 (now), so they decided to have maintenance today.  That will continue until 20:00

Several things were accomplished during the extended access.
1. ShowerMax was installed and cabled.
2. Bravho was modified to control ShowerMax
3. DAQ was restored from its "wierd" state after the power dip.
4. TRMRS was restored and MRS1440 communication was restored.
5. D1 is in a Stby-Fault state and cannot be controlled.  Called MCR and they said they would send someone once things calm down over there.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Tue Feb 1 16:15:59 2005

This archive was generated by hypermail 2.1.8 : Tue Feb 01 2005 - 16:16:24 EST